# A Novel Federated Learning-Based IDS for Enhancing UAVs Privacy and Security

### v2

@@ II Related Work
[22] 2022 FANETs UAV-IDS-2020 [23] attacks for unauthorized access in WLAN ML + data collected from real UAVs were used.
- dataset is not suitable for FANETs
[10] 2023 UAVs UAV dataset GPS Spoofing ML + various ML-based algorithms are tuned and evaluated
(a real UAV was used)
[18] 2023 FANETs a FANET dataset Sybil ML + a new dataset is created
(100 nodes, 3D INET’s MassMobility) - it is evaluated on a small area with nodes having low speeds (10-20 m/s)
@@
[20] 2023 FANETs a FANET dataset Time Delay ML + a new dataset is created by using 4 different routing protocols
@@
[29] 2019 FANETs a FANET dataset collected from nodes Spoofing, Jamming DL + it is claimed to be lightweight
@@
[24] 2021 FANETs KDDCup’99 [25] DoS, DDoS, Probe, R2L, U2R DL + both local and central IDSs participate in detection
NSL-KDD [26] Fuzzer, Analysis, Backdoors, Exploits - the datasets are not suitable for FANETs
@@
[27] 2021 FANETs CICIDS2017 [28] Brute force, DoS, DDoS DL + offline-learning is proposed, where UAVs can update models during charging
BotNet, Web Attack, Infiltration - the dataset is not suitable for FANETs
[30] 2019 FANETs CRAWDAD [34] for VANETs Jamming FL + a client group prioritization technique is proposed
a FANET dataset (6 nodes, 3D GM Mobility Model) - the effect of mobility on client group prioritization is not discussed
@@
[42] 2020 IoT NSL-KDD [26] DoS, U2R, R2L, Probing attacks for training FL + data distribution is explored for different types of attacks
@@
[41] 2021 IoT Bot-IoT [48] DoS, DDoS, Theft, Reconnaissance FL + different learning algorithms are evaluated
MQTTset [49] Specific Attacks against MQTT + the dataset is collected from heterogeneous devices
@@
[38] 2022 IoT ToN_IoT[39] Backdoor,Dos,DDos,Injection,MITM FL + realistic dataset with consideration of non-independent and identically distributed data
Password,Ransomware,Scanning,XSS - resource consumption is not examined
[32] 2023 UAV swarm WSN-DS [47] Blackhole, Grayhole, Flooding ML + various ML-based algorithms are evaluated
UAV attack dataset [35] Ping DoS - a significant percentage of the normal class is classified as grayhole
@@
Our 2024 FANETs a FANET dataset Sinkhole, Blackhole, Flooding FL + a new dataset is created
Study (50 nodes, 3D GM Mobility Model, + extensive simulations are carried out
@@
high speed: 720km/h) - communication cost during training
III Background
@@
III-A Federated Learning
The utilization and processing of data collected from distributed devices pose challenges for traditional centralized IDSs. Firstly, there is the violation of privacy concern related to data sharing. In the conventional approach, the central server has access to all private client data to train a global model, which is then shared with all clients [50]. However, this raises privacy issues, as the data may contain private information that clients prefer not to share with third parties. Secondly, in the context of FANETs characterized by high node speeds and frequent link breakages, relying on communication among IDS clients may not be preferable due to potential disruptions and unreliability in data transmission. This exacerbates the challenges posed by dealing with large volumes of data in a distributed environment, resulting in increased computational and communication costs, ultimately causing latency in a central IDS. Lastly, energy consumption becomes a significant concern for resource-constrained devices, and both clients and central systems might require high energy consumption in order to obtain, store, process the local and the aggregated data respectively. Solutions that consider energy consumption must be designed for FANETs, particularly due to the reliance of mini UAVs on low-capacity batteries for power [51].
Figure 1: Conceptual Schema of Federated Learning
@@ III-A Federated Learning
  • Aggregation: The updates from all participating devices or servers are sent back to the central server and combined to produce a new version of the model.
@@
  • Iteration: The process of training, updating, and aggregating is repeated until the model reaches an acceptable level of accuracy.
@@
FL is a promising approach for our study, driven by several key advantages. Most notably, FL significantly enhances efficiency by mitigating the need to transmit large amounts of data to a central location for processing, resulting in saving in both time and bandwidth. This reduction in the cost of data storage and processing renders FL a cost-effective solution, particularly beneficial for resource-constrained devices and environments.
@@
Moreover, its inherent scalability and flexibility make FL adaptable to a wide variety of AI tasks. Notably, FL inherently prioritizes privacy, a paramount concern in today’s data-driven landscape. Moreover, FL’s capacity to handle missing data and adapt to data drifts (common occurrences in real-world applications) ensures the development of robust models [52]. Consequently, our decision to utilize FL aligns seamlessly with our overarching aim: to harness the power of ML while upholding data privacy and security. This proposition is particularly appealing for IDSs seeking effective, privacy-preserving, and distributed AI solutions.
III-B Ad hoc On-Demand Distance Vector (AODV)
@@
The AODV routing protocol is a reactive routing protocol initially designed for MANETs. Unlike proactive routing protocols that maintain a complete network topology, AODV creates routes only when needed by a node to transmit data to another node. When a node intents to transmit data to another node beyond its immediate range, it initiates a route discovery process by broadcasting a route request (RREQ) packet. This RREQ packet propagates through the network until it reaches either the destination node or a node that has a sufficiently fresh route to the destination. Subsequently, the route is established by sending a route reply (RREP) packet back to the source node. During this process, when the destination node or an intermediate node processing an active route to the destination in its routing table receive a RREQ packet, it sends a unicast RREP packet to the source node. The source node then selects the shortest and most up-to-date route based on the minimum hop count and the highest destination sequence number, respectively. This route selection ensures the transmission of data through the shortest and most up-to-date route available.
@@
AODV also incorporates a route maintenance mechanism to effectively handle link failures and route breaks. This process begins with the transmission of a Route Error (RERR) message. Upon detecting a link break or node failure, the node generates and sends the RERR message to nodes possessing a route to the unreachable destination. This action promptly invalidates the affected routes. Subsequently, the affected node initiates a new route discovery process to find an alternative path to the destination. AODV is highly regarded in the context of FANETs due to its low overhead and its rapid adaptability to network changes. AODV has been used in a variety of FANETs applications [53].
III-C Attacks
@@ III-C1 Sinkhole Attack
III-C2 Blackhole Attack
A blackhole attack combines elements of sinkhole and dropping attacks, enabling the attacker to selectively discard packets destined for a specific destination. The attacker can choose to drop all received packets to disrupt network communication or employ random dropping to evade detection. Similar to a sinkhole attack, the blackhole attacker initially advertises itself as having the best route to the desired destination, attracting network traffic towards it. Subsequently, additional attacks such as data modification and packet dropping may be executed. As a result, a blackhole attack can severely disrupt network communication and lead to increased energy consumption.
III-C3 Flooding Attack
@@ IV The Proposed Approach
IV-A Network and Attack Settings
We conducted the three distinct attacks, introduced earlier, namely sinkhole, blackhole and ad hoc flooding, against the AODV routing protocol within FANETs, utilizing the Ns-3 simulator [33]. Each network consisted of 50 mobile nodes to simulate the multi-hop characteristics of ad hoc networks. Additionally, an immobile Ground Base Station (GBS) node was positioned at the simulation center due to its utility in various applications, including disaster response, agricultural monitoring, and military operations [55, 56]. The GBS serves as a central hub for receiving, processing, and analyzing data collected by the UAVs. With this setting, our scenario simulates both the Air-to-Air (A2A) communication channel and the UAV to GBS link as Air-to-Ground (A2G) channel communication[43, 57].
The 3D natural flight of UAVs was emulated with the 3D Gauss Markow (GM) mobility model to provide a realistic approach [58]. This model mimics real-world scenarios, producing smoother turns and predictable movement paths rather than entirely random motion. Here, high-speed drones that can be used in military applications [59] are simulated. Within the network, 10 source nodes and 10 destination nodes communicate randomly, with any remaining nodes potentially serving as relay nodes. Data collection and transmission to the GBS are the responsibilities of the destination nodes, commencing at the 10th second and continuing every second until the end of the simulation.
Attacker nodes were randomly selected from non-source or non-destination nodes. For each attacker ratio, ranging from 5% to 25%, attacker nodes were chosen randomly. These same attacker nodes were used consistently in each attack type. The network simulation parameters are listed in Table II. These parameters were carefully chosen to align with real-life scenarios, considering factors such as 3D natural flight patterns, realistic speeds, area selection, and an adaptable routing protocol for dynamic FANET movements. More details on parameter selection can be found in our previous study [60].
TABLE II: Network Simulation Parameters
@@ IV-A Network and Attack Settings
TABLE III: Dataset Information
@@
Blackhole Attack: The blackhole attack combines elements of the sinkhole and dropping attacks. Initially, the attacker establishes a route, attracting traffic as described in the sinkhole attack. Subsequently, it indiscriminately drops every packet it receives, thus disrupting communication between endpoints. However, the attacker’s position is critical in dropping attacks. If located on a passive route, it has no impact on network performance. Due to this, and the need for knowledge from other layers to differentiate it from dropping caused by wireless links, dropping attacks alone were not used in the simulations.
@@
Flooding Attack: In this scenario, an attacker node periodically sends RREQ packets to randomly selected destinations within the network. The chosen destination node receives multiple (=10) sequential RREQ broadcast messages, with this process being repeated every 3 seconds throughout the simulation.
@@
To ensure a comprehensive evaluation, we created ten distinct network topologies with diverse network and mobility patterns, initially without any attack implementations. Subsequently, we introduced malicious nodes into these networks, enabling us to simulate a wide range of attack scenarios. We conducted 10 simulations for each attack, each with a different attacker ratio, ranging from 5% to 25%. Additionally, we ran 10 simulations without any attacks, totaling 160 simulations. Data were collected from each simulation over a period of 1800 seconds.
@@
In our data collection process, we gathered features introduced in [61] from each node. The features extracted encompass a range of characteristics, some related to mobility, while others related to AODV control messages and data packets, as detailed in [61]. For instance, there are features that provide insights into the frequency of forwarding, sending and reception of routing protocol control packets (RREQ, RREP, and RERR). Additionally, there are features that give information about changes in the routing table. These features are collected periodically from each node, with a feature collection period set to 5 seconds. The dataset information is summarized in Table III. Please note that while the frequency of data collection positively affects the performance of the system, it also increases resource consumption. Further exploration of trade-offs in the selection of this parameter can be conducted in future studies.
IV-B Federated IDS
@@
The study introduces a Federated Learning-based Intrusion Detection System (FL-IDS) for detecting network attacks in FANETs and compares it to traditional methods. Three types of IDS are considered: Central IDS (C-IDS), Local IDS (L-IDS), and FL-IDS. The impact of both traditional and federated learning methods on IDS is assessed in terms of accuracy in this section. They are compared using different criteria including effectiveness, efficiency, communication, privacy and security in the subsequent section. All IDS models utilize Deep Neural Networks (DNN) and Convolutional Neural Networks (CNN). To facilitate model training and testing, the dataset is divided into 80% for training and 20% for testing, following the application of a standard scaler. The details of each IDS are given in the following:
@@
  • L-IDS: In this approach, each node is individually trained with its own local data to create its own model. Each local model is then deployed to all UAVs. Please note that the presented evaluation result is the average of all L-IDS results.
@@
  • FL-IDS: In the proposed method, each UAV is initially assigned a default local model to incorporate client-specific data in the IDS. Each client/node trains the default model with their local data, creating new weights. After this process, only the fresh model’s weights are shared with the central server located at the GBS. The server aggregates weights from each client, updating the global model with the average of these weights using the FedAvg method [36]. The updated global model’s weights are then sent back to each client for further training iterations, and these processes continue until the defined epoch.
@@
These approaches have a number of parameters that need to be tuned for optimal performance. To find the best parameters for the models, the Grid Search method is employed. This method exhaustively searches a predefined set of hyperparameters and selects the combination that produces the best performance. The best parameters selected for optimum performance are shown in Table IV. The performance of models is evaluated on the test set using the following metrics: accuracy, detection rate (DR) and false positive rate (FPR). The final results are obtained by averaging the results of the method applied to different topologies for each IDS.
@@
In summary, our study involves three different types of IDS, and we apply CNN and DNN algorithms to the dataset to create models for these IDSs. As mentioned earlier, the Federated IDS (FL-IDS) addresses the disadvantages of both C-IDS and L-IDS, such as increased communication costs and privacy violations. This approach allows the central server to aggregate model weights without directly accessing raw data from individual nodes. By sharing only the model weights, we effectively address privacy concerns while still enabling collaborative analysis and intrusion detection in a secure manner. This approach is particularly well-suited for distributed systems like FANETs. It is worth noting that most studies in this field [41] do not reflect real-world client-specific data distribution due to a lack of FL-specific datasets. However, in our study, we collected client-based data, providing a more accurate reflection of real-world client-specific data distribution.
TABLE IV: Deep Learning Parameters
@@
V Experimental Results
In this section, we present and analyze the outcomes of our experiments, providing a comprehensive evaluation of the proposed FL-IDS. The performance of FL-IDS is compared with traditional intrusion detection methods, namely Central IDS (C-IDS) and Local IDS (L-IDS), across various attack scenarios, including sinkhole, blackhole, and flooding attacks, under different attacker ratios in the FANET. Additionally, we introduce and discuss the effectiveness of the Bias Towards Specific Clients (BTSC) method, which further refines FL-IDS performance. The experimental results offer valuable insights into the strengths and capabilities of FL-IDS in addressing the security challenges posed by dynamic and decentralized UAV networks.
V-A Classifier: DNN vs CNN
In this subsection, the performance of DNN and CNN algorithms is compared for all simulated attack types. The results are shown in Tables V, VI, VII for sinkhole, blackhole, and flooding attacks, respectively. Across all IDSs, CNN outperforms DNN. The same trend is observed for the proposed FL-based approach. In FL-IDS, while CNN is better than DNN across all attacker ratios, the gap between the two algorithms widens as the attacker ratio increases. Since CNN consistently outperforms DNN, we will present only the CNN results in the subsequent sections.
TABLE V: Accuracy of DNN and CNN in Detecting Sinkhole Attack

### v3

@@ 2 Related Work
[22] 2022 FANETs UAV-IDS-2020 [23] attacks for unauthorized access in WLAN ML Accuracy, Precision, TPR +Data collected from real UAVs were used.
Recall, F1-score, FPR -Dataset is not suitable for FANETs
[10] 2023 UAVs UAV dataset GPS Spoofing ML Accuracy, Precision, TPR +Various ML-based algorithms are tuned and evaluated
(a real UAV was used) Recall, F1-score, FPR
[18] 2023 FANETs a FANET dataset Sybil ML Accuracy +A new dataset is created
(100 nodes, 3D INET’s MassMobility) TPR -It is evaluated on a small area with nodes having low speeds (10-20 m/s)
@@
[20] 2023 FANETs a FANET dataset Time Delay ML Accuracy +A new dataset is created by using 4 different routing protocols
@@
[29] 2019 FANETs a FANET dataset collected from nodes Spoofing, Jamming DL Accuracy +It is claimed to be lightweight
@@
[24] 2021 FANETs KDDCup’99 [25] DoS, DDoS, Probe, R2L, U2R DL Accuracy, Precision +Both local and central IDSs participate in detection
NSL-KDD [26] Fuzzer, Analysis, Backdoors, Exploits TPR, F1-score -The datasets are not suitable for FANETs
@@
[27] 2021 FANETs CICIDS2017 [28] Brute force, DoS, DDoS DL Accuracy, Precision +Offline-learning is proposed, where UAVs can update models during charging
BotNet, Web Attack, Infiltration Recall, F1-score -The dataset is not suitable for FANETs
[30] 2019 FANETs CRAWDAD [34] for VANETs Jamming FL Accuracy +A client group prioritization technique is proposed
a FANET dataset (6 nodes, 3D GM Mobility Model) Running Time -The effect of mobility on client group prioritization is not discussed
@@
[44] 2020 IoT NSL-KDD [26] DoS, U2R, R2L, Probing attacks for training FL Accuracy +Data distribution is explored for different types of attacks
@@
[43] 2021 IoT Bot-IoT [50] DoS, DDoS, Theft, Reconnaissance FL Accuracy, Precision +Different learning algorithms are evaluated
MQTTset [51] Specific Attacks against MQTT Recall, F1-score +The dataset is collected from heterogeneous devices
@@
[38] 2022 IoT ToN_IoT[39] Backdoor,Dos,DDos,Injection,MITM FL Accuracy, Precision +Realistic dataset with consideration of non-independent and identically distributed data
Password,Ransomware,Scanning,XSS Recall, F1-score, FPR -Resource consumption is not examined
[41] 2023 IoT UNSW-NB15 [48] Fuzzers, Analysis, Backdoors FL Accuracy, Precision +Combines FL with Blockchain to ensure privacy, security, and trust
DoS, Exploits, Generic Blockchain Recall, F1-score +Uses SMPC for secure aggregation, preserving data privacy
Reconnaissance, Shellcode and Worms Training Time -Increased computational and communication overhead
-Blockchain implementation can be complex and resource-intensive
[32] 2023 UAV swarm WSN-DS [49] Blackhole, Grayhole, Flooding ML F1-score, AUC +Various ML-based algorithms are evaluated
UAV attack dataset [35] Ping DoS Training Time -A significant percentage of the normal class is classified as grayhole
@@
Our 2025 FANETs a FANET dataset Sinkhole, Blackhole, Flooding FL Accuracy, Recall +A new dataset is created
Study (50 nodes, 3D GM Mobility Model, FPR +Extensive simulations are carried out
@@
average speed: 100 m/s) -Communication cost during training
3 Background
@@
3.1 Federated Learning
The utilization and processing of the data collected from distributed devices pose challenges for traditional centralized IDSs. Firstly, there is the violation of privacy concern related to data sharing. In the conventional approach, the central server has access to all private client data to train a global model, which is then shared with all clients [52]. However, this raises privacy issues, as the data may contain private information that clients prefer not to share with third parties. Secondly, in the context of FANETs characterized by high node speeds and frequent link breakages, relying on communication among IDS clients may not be preferable due to potential disruptions and unreliability in data transmission. This exacerbates the challenges posed by dealing with large volumes of data in a distributed environment, resulting in increased computational and communication costs, ultimately causing latency in a central IDS. Lastly, energy consumption becomes a significant concern for resource-constrained devices, and both clients and central systems might require high energy consumption in order to obtain, store, process the local and the aggregated data respectively. Solutions that consider energy consumption must be designed for FANETs, particularly due to the dependence of mini UAVs on low-capacity batteries for power [53].
Figure 1: Architectural Overview of of Federated Learning
@@ 3.1 Federated Learning
    Aggregation: Updates from all participating devices or servers are sent back to the central server and combined to produce a new version of the model.
  • 4.
@@
    Iteration: The training, updating and aggregating process is repeated until the model reaches an acceptable level of accuracy.
@@
FL is a promising approach for our study, driven by several key advantages. In particular, FL significantly enhances efficiency by mitigating the need to transmit large amounts of data to a central location for processing, resulting in savings in both time and bandwidth. This reduction in the cost of data storage and processing renders FL a cost-effective solution, particularly beneficial for resource-constrained devices and environments.
@@
Moreover, its inherent scalability and flexibility make FL adaptable to a wide range of AI tasks. Notably, FL prioritizes privacy, a paramount concern in today’s data-driven landscape. Additionally, FL’s ability to handle missing data and adapt to data drifts (common occurrences in real-world applications) ensures the development of robust models [54]. Consequently, our decision to utilize FL aligns seamlessly with our overarching aim: to harness the power of ML while upholding data privacy and security. This proposition is particularly appealing for IDSs seeking effective, privacy-preserving, and distributed AI solutions.
3.2 Ad hoc On-Demand Distance Vector (AODV)
@@
The AODV routing protocol is a reactive routing protocol originally designed for MANETs. Unlike proactive routing protocols that maintain a complete network topology, AODV creates routes only when a node needs to transmit data to another node. When a node intends to transmit data to another node beyond its immediate range, it initiates a route discovery process by broadcasting a route request (RREQ) packet. This RREQ packet propagates through the network until it reaches either the destination node or a node that has a sufficiently fresh route to the destination. Subsequently, the route is established by sending a route reply (RREP) packet back to the source node. During this process, when the destination node or an intermediate node processing an active route to the destination in its routing table receives an RREQ packet, it sends a unicast RREP packet to the source node. The source node then selects the shortest and most up-to-date route based on the minimum hop count and the highest destination sequence number, respectively. This route selection ensures the transmission of data through the shortest and most up-to-date route available.
@@
AODV also incorporates a route maintenance mechanism to effectively handle link failures and route breakdowns. This process begins with the transmission of a Route Error (RERR) message. Upon detecting a link break or node failure, the node generates and sends the RERR message to the nodes that possess a route to the unreachable destination. This action promptly invalidates the affected routes. Subsequently, the affected node initiates a new route discovery process to find an alternative path to the destination. AODV is highly regarded in the context of FANETs due to its low overhead and its rapid adaptability to network changes.
3.3 Attacks
@@ 3.3.1 Sinkhole Attack
3.3.2 Blackhole Attack
A blackhole attack combines elements of sinkhole and dropping attacks, enabling the attacker to selectively discard packets destined for a specific destination. The attacker can choose to drop all received packets to disrupt network communication or employ random dropping to evade detection. Similarly to a sinkhole attack, the blackhole attacker initially advertises itself as having the best route to the desired destination, attracting network traffic towards it. Subsequently, additional attacks, such as data modification and packet dropping may be executed. As a result, a blackhole attack can severely disrupt network communication and lead to increased energy consumption.
3.3.3 Flooding Attack
@@ 4 The Proposed Approach
4.1 Network and Attack Settings
We conducted the three distinct attacks, introduced earlier, namely sinkhole, blackhole and ad hoc flooding, against the AODV routing protocol within FANETs, utilizing the Ns-3 simulator [33]. Each network consisted of 50 mobile nodes to simulate the multi-hop characteristics of ad hoc networks. Furthermore, an immobile Ground Base Station (GBS) node was positioned at the simulation center due to its utility in various applications, including disaster response, agricultural monitoring, and military operations [56, 57]. The GBS serves as a central hub for receiving, processing, and analyzing data collected by the UAVs. With this setting, our scenario simulates both the Air-to-Air (A2A) communication channel and the UAV to GBS link as Air-to-Ground (A2G) channel communication[45, 58].
The 3D natural flight of UAVs was emulated with the 3D Gauss Markov (GM) mobility model to provide a realistic approach [59]. This model mimics real-world scenarios, producing smoother turns and predictable movement paths instead of entirely random motion. Within the network, 10 source nodes and 10 destination nodes communicate randomly, with any remaining nodes potentially serving as relay nodes. Data collection and transmission to the GBS are the responsibilities of the destination nodes, beginning at the 10th second and continuing every second until the end of the simulation.
The attack nodes were randomly selected from the non-source or non-destination nodes. For each attacker ratio, ranging from 5% to 25%, the attacker nodes were randomly chosen. These same attacker nodes were used consistently in each type of attack. The network simulation parameters are listed in Table 2. These parameters were carefully chosen to align with real-life scenarios, considering factors such as 3D natural flight patterns, area selection, and an adaptable routing protocol for dynamic FANET movements. More details on parameter selection can be found in our previous study [60].
Table 2: Network Simulation Parameters
@@ 4.1 Network and Attack Settings
Table 3: Dataset Information
@@
Blackhole Attack: The blackhole attack combines elements of the sinkhole and dropping attacks. Initially, the attacker establishes a route, attracting traffic as described in the sinkhole attack. Subsequently, it indiscriminately drops every packet it receives, thus disrupting communication between endpoints. However, the attacker’s position is critical in dropping attacks. If it is located on a passive route, it has no impact on network performance. Due to this and the need for knowledge from other layers to differentiate it from dropping caused by wireless links, dropping attacks alone were not used in the simulations.
@@
Flooding Attack: In this scenario, an attacker node periodically sends RREQ packets to randomly selected destinations within the network. The chosen destination node receives multiple sequential (= 10) RREQ broadcast messages, with this process repeated every 3 seconds throughout the simulation.
@@
To ensure a comprehensive evaluation, we created ten different network topologies with diverse network and mobility patterns, initially without any attack implementations. Subsequently, we introduced malicious nodes into these networks, enabling us to simulate a wide range of attack scenarios. We conducted 10 simulations for each attack, each with a different attacker ratio, ranging from 5% to 25%. Additionally, we ran 10 simulations without any attacks, totaling 160 simulations. Data were collected from each simulation over a period of 1800 seconds.
4.2 Federated IDS
@@
The study introduces a Federated Learning-based Intrusion Detection System (FL-IDS) to detect network attacks in FANETs and compares it to traditional methods. Figure 2 illustrates the main steps of the proposed approach. The first step is data collection, during which we collected all the features introduced in [61] from each node. The extracted features cover a wide range of characteristics, including those related to mobility, as well as those associated with AODV control messages and data packets, as described in [61]. For example, certain features provide information on the frequency of forwarding, sending, and receiving routing protocol control packets (RREQ, RREP, and RERR). Additionally, some features reflect changes in the routing table, which can be attributed to mobility. Furthermore, features such as variations in the number of neighbors offer direct indications of mobility patterns. These features were collected from each node at 5-second intervals. The dataset information is summarized in Table 3. Please note that while the frequency of data collection positively affects the performance of the system, it also increases resource consumption. Further exploration of trade-offs in the selection of this parameter can be conducted in future studies.
@@
To facilitate model training and testing, the dataset was split into 80% for training and 20% for testing, with a standard scaler applied. All IDS models employ Deep Neural Networks (DNN) and Convolutional Neural Networks (CNN). Three types of IDS are considered: Central IDS (C-IDS), Local IDS (L-IDS), and FL-IDS. The performance of traditional and federated learning methods on IDS is assessed in terms of accuracy and is compared across criteria such as effectiveness, efficiency, communication, privacy, and security in subsequent sections.
The following subsections detail the specifics of each IDS.
  • 1.
@@
    L-IDS: In this approach, each node is individually trained with its own local data to create its own model. Then each local model is deployed to all UAVs. Please note that the evaluation result presented is the average of all L-IDS results.
  • 3.
@@
    FL-IDS: In the proposed method, each UAV is initially assigned a default local model to incorporate client-specific data in the IDS. Each client/node trains the default model with their local data, creating new weights. After this process, only the fresh model’s weights are shared with the central server located at the GBS. The server aggregates the weights of each client, updating the global model with the average of these weights using the FedAvg method [36]. The updated global model’s weights are then sent back to each client for further training iterations, and these processes continue until the defined epoch.
    To address the dynamic and heterogeneous nature of FANETs, we employed a methodology called Bias Towards Specific Clients (BTSC) to enhance the performance of our FL-IDS. BTSC prioritizes clients who demonstrate superior attack detection capabilities by assigning higher weights to their model updates during the federated aggregation process. Clients are ranked on the basis of their detection performance metrics, ensuring that the most effective models contribute more significantly to the global model. The top 20% of clients, identified from previous experiments based on their high detection accuracy, are selected for aggregation.
    To mitigate the risk of overfitting to these high-performing clients, the global aggregated model is evaluated on a comprehensive data set that includes data from all clients, both contributing and non-contributing. This evaluation process ensures that the model is effectively generalized across diverse data distributions, maintaining robust detection performance without bias toward specific clients. Our results confirm that the approach improves the effectiveness of the global model while preserving its generalizability.
@@
These approaches have a number of parameters that need to be tuned for optimal performance. To find the best parameters for the models, the Grid Search method is employed. This method exhaustively searches for a predefined set of hyperparameters and selects the combination that produces the best performance. The best parameters selected for optimal performance are shown in Table 4.
The performance of models is evaluated on the test set using the following metrics: accuracy, detection rate (DR) and false positive rate (FPR). The final results are obtained by averaging the results of the method applied to different topologies for each IDS.
@@
This novel, distributed approach is specifically designed to align with the unique characteristics of FANETs, such as their high mobility, decentralized architecture, and dynamic network topologies. Using federated learning, FL-IDS reduces communication costs and preserves privacy by sharing only model weights instead of raw data. This allows the central server to aggregate model weights without directly accessing raw data from individual nodes, addressing the disadvantages of both C-IDS and L-IDS, such as increased communication costs and privacy violations. By sharing only the model weights, we effectively mitigate privacy concerns while enabling collaborative analysis and intrusion detection in a secure manner. This approach is particularly well-suited for distributed systems like FANETs. Our method is applied to routing attacks, using a dataset specifically tailored to FANETs that includes features like 3D mobility patterns and realistic data collection, ensuring a practical and accurate evaluation of the system. It is worth noting that most studies in this field [43] do not reflect the distribution of client-specific data in the real world due to the lack of FL-specific datasets.
Table 4: Deep Learning Parameters
@@
5 Experimental Results
In this section, we present and analyze the results of our experiments, providing a comprehensive evaluation of the proposed FL-IDS. The performance of FL-IDS is compared with traditional intrusion detection methods, namely Central IDS (C-IDS) and Local IDS (L-IDS), across various attack scenarios, including sinkhole, blackhole, and flooding attacks, under different attacker ratios in the FANET. In addition, we introduce and discuss the effectiveness of the Bias Towards Specific Clients (BTSC) method, which further refines FL-IDS performance. The experimental results offer valuable insights into the strengths and capabilities of FL-IDS in addressing the security challenges posed by dynamic and decentralized UAV networks.
5.1 Classifier: DNN vs CNN
In this subsection, the performance of the DNN and CNN algorithms is compared for all types of simulated attacks. The results are shown in Tables 5, 6, 7 for sinkhole, blackhole, and flooding attacks, respectively. Across all IDSs, CNN outperforms DNN. The same trend is observed for the proposed FL-based approach. In FL-IDS, while CNN is better than DNN across all attacker ratios, the gap between the two algorithms widens as the attacker ratio increases. Since CNN consistently outperforms DNN, we will present only the CNN results in the subsequent sections.
Table 5: Accuracy of DNN and CNN in Detecting Sinkhole Attack